### v3

# Network Analytics for Anti-Money Laundering – A Systematic Literature Review and Experimental Evaluation

Abstract
Money laundering presents a pervasive challenge, burdening society by financing illegal activities. The use of network information is increasingly being explored to more effectively combat money laundering, given it involves connected parties. This led to a surge in research on network analytics (NA) for anti-money laundering (AML). The literature on NA for AML is, however, fragmented and a comprehensive overview of existing work is missing. This results in limited understanding of the methods to apply and their comparative detection power. Therefore, this paper presents an extensive and unique literature review, based on 97 papers from Web of Science and Scopus, resulting in a taxonomy following a recently proposed fraud analytics framework. We conclude that most research relies on expert-based rules and manual features, while deep learning methods have been gaining traction. This paper also presents a comprehensive framework to evaluate and compare the performance of prominent NA methods in a standardized setup. We apply it on two publicly available data sets, comparing manual feature engineering, random walk-based, and deep learning methods. We conclude that (1) network analytics increases the predictive power, but caution is needed when applying GNNs based on the class imbalance and network topology, and that (2) care should be taken with open-source data as this can give overly optimistic results. The open-source implementation facilitates researchers and practitioners to extend upon the results and experiment on proprietary data, promoting a standardized approach for the analysis and evaluation of network analytics for AML.
Keywords Fraud Analytics  $\cdot$ Anti-Money Laundering  $\cdot$ Network Analytics  $\cdot$ Literature Review
1 Introduction
Money laundering is the process of concealing illegally obtained funds by passing them through a complex sequence of transactions. Laundered money can be used in the legitimate financial system to fund and facilitate criminal activities, such as drug trafficking, corruption, and terrorism [77]. Around $2\%$ to $5\%$ of global GDP is laundered worldwide, amounting yearly to USD 2 trillion, as estimated by the United Nations Office on Drugs and Crime (UNODC) [130].
Money laundering differs from other forms of financial fraud, especially credit card and payment fraud, as it is often done over a longer period and only becomes visible when looking at multiple transactions. It is more actively concealed, since sustained laundering is needed to process continuous streams of money. Credit card fraud, on the other hand, is much more about getting as much money out as fast as possible. Hence, the specific characteristics of money laundering merit dedicated research attention.
Since, this illegally obtained money needs to enter the financial system at some point [77], legislative bodies have put stringent rules on financial institutions [15], requiring them to have safeguards and to report suspicious transactions/activities to the Financial Intelligence Unit (FIU). Despite record fines and mounting pressure by regulators, money laundering represents a growing problem [140]. Anti-money laundering practices are largely ineffective [35], with Europol estimating only 1.1% of criminal profits in the EU being confiscated [38].
Money laundering generally requires three main steps [130, 77]. The first is placement, where dirty money is integrated into the financial system. The second is layering, where money is transferred and mixed to obscure its illegal origin. The third is integration, where the laundered money is integrated into wider circulation by doing seemingly legitimate purchases.
Anti-money laundering practices are focused on placement and layering, as after integration via legitimate purchases, the money is in essence successfully laundered. Know-Your-Customer (KYC) protocols screen novel clients when they want to do business with the bank. Afterwards, transactions are continuously monitored to see if a client is involved in money laundering.
Anti-money laundering practices are often a combination of rule-based and risk-based systems [57, 123]. Rule-based systems consist of a collection of different IF-THEN rules. These are defined based on regulatory requirements and expert knowledge [110]. In practice, rule-based systems are still widely used [41, 22, 110]. The main drawback is that many rules are needed as money laundering is complex, which are hard to maintain. Risk-based systems apply more sophisticated methods. Past literature has shown that popular data analytics and machine learning methods have been applied for anti-money laundering [100, 57, 15, 76]. However, Le Khac and Kechadi [76] claim that most commercial tools still rely on more simple statistics to find outliers.
Placement, layering and integration involve multiple parties, and individual transactions often appear normal [15]. Therefore, network analytics is required to uncover the specific money laundering characteristics [15, 10]. The need for network analytics was already mentioned three decades ago by Senator et al. [116], where the authors used visualisations of relations to support experts investing suspicious cases. Although Ngai et al. [100] said that there was still a lack of research at that time, more attention has been given to money laundering and network analytics in the past years.
The increasing attention has resulted in a growing body of literature covering network analytics for anti-money laundering. The swift expansion of the field has simultaneously resulted in a lack of a comprehensive overview. Most literature reviews on financial fraud barely discuss anti-money laundering using network analytics [100, 75, 42, 56]. Therefore, as a first step, this paper aims to supplement the literature by providing a systematic literature review of network analytics (NA) for anti-money laundering. To this end, we constructed queries to identify relevant papers from Web of Science (WOS) and Scopus, resulting in 97 papers.
The papers are analysed according to different dimensions. First, we look at the evolution of the number of published papers over time. Second, each paper is classified according to seven categories; publication data, the method type, the modelling method, the evaluation metrics used, the main objective, the kind of data, and network characteristics. We identify an increasing focus on crypto currencies. Our work compares the differences between those papers and the research covering classic financial transactions. Third, the top-cited work is reviewed via the framework introduced by Bockel-Rickermann et al. [14].
Our literature review and subsequent analysis reveals that multiple supervised, unsupervised and visualisation methods have been developed and applied. Nevertheless, due to the literature being fragmented and these methods mostly addressing only a part of the AML problem, there is limited insight into how different methods compare and which work best. Furthermore, most state-of-the-art (SOTA) network analytics methods have not yet found main-stream adoption for anti-money laundering. Therefore, this work provides a comprehensive experiment, evaluating the most prominent methods.
In summary, our main contributions are four-fold:
  • We provide an extensive and critical systematic literature review on network analytics for anti-money laundering, to gain both comprehensive and deep insights, and provide direction for future research.
  • We benchmark a wide range of state-of-the-art network analytics methods in the context of anti-money laundering, comparing the performance of these key methods on two open-source data sets: the Elliptic [37, 139] and the IBM Transactions for AML [6, 36] data set.
  • Based on our findings, we provide insights into the specific challenges of network methods based on the topology of the network and the data generating process.
  • We implement the methods in a uniform manner and facilitate replication of the presented results by providing public access to the code at https://github.com/B-Deprez/AML_Network, aiming to promote a standardized approach towards the analysis and adoption of NA for AML and to encourage further research.
This paper is structured as follows. We present the methodology for the structured literature review in Section 2, with Section 3 presenting the results of the review. Section 4 describes the set-up of the empirical evaluation, with the results presented in Section 5. Finally, overall conclusions and directions for future work are presented in Section 6.
2 Literature Review
2.1 Methodology
We conducted a systematic literature search using the queries “graph analy*” AND “money launder*” and “network analy*” AND “money launder*” on WoS and Scopus, chosen for their high completeness and research credibility [97, 14]. The search is limited to English-language papers published before 2023, excluded recent publications to ensure adequate time for community evaluation.
The selection process included a title and abstract scan, followed by a full-text review. Papers were excluded if they focused on legal aspects or lacked substantial application of network analytics to anti-money laundering (AML). The final set of papers was classified according to seven criteria: publication data, method type, modelling approach, evaluation metrics, main objectives, data type, and network characteristics.
Our review provides both a broad analysis of trends across all selected papers and an in-depth focus on top-cited works (top $10\%$ annually and overall, per Google Scholar). This dual approach highlights general trends and the state-of-the-art (SOTA) in the field. Finally, we examine relevant review papers to contextualize our findings and demonstrate how our work extends existing literature.
Table 1: The (sub-)categories with explanation for the classification of the literature.
Category Sub-Category Definition
Publication Data Title The full title of the paper
Journal/Conference Where the paper is published
Year The year of publication
Review Paper Indicating if it is a review paper
Bitcoin/Crypto Indicating if it deals with crypto currency
Method Type Supervised Application of supervised method
Unsupervised Application of unsupervised method
Semi-Supervised Application of semi-supervised method
Mixed Methods presented across different learning methods
Visualisation Application of visualisation method
Modelling Method Rule-based Application of specific rules or cut-offs
Manual features Network features having a exact definition
Walk-based Network analysis based on (random) walks
Shallow representation Embeddings typically through matrix factorization or random walks, without employing deep neural networks
Deep representation Embeddings based on deep neural networks, e.g., GNNs
Correlation-based Analysis of correlation of network features with target
Logistic regression Application of logistic regression
Tree-based Application of tree-based methods
SVM-based Application of support vector machines
Neural networks Learning using neural networks
Anomaly detection Application of anomaly or outlier detection
Clustering Features based on community detection or clustering of features
Evaluation Metric Accuracy Accuracy
Precision The precision
Recall The recall
F1 The F1 or micro-F1 score
TPR The true positive rate
FPR The false positive rate
AUC-ROC The area under the ROC curve
AUC-PR The area under the precision-recall curve
Time The execution time
Objective Client classification Detection of suspicious entities
Transaction classification Detection of suspicious transactions
Community detection Detection of suspicious groups of clients
Flow/Chain detection Detection of suspicious combination of transactions
Data Proprietary Usage of confidential data
Open-source Usage of freely available data
Synthetic Usage of synthetically generated data
Network Multiple Usage of multiple networks
Inverse/undirected Additional use of network with directions removed or reverted
Web of ScienceScopusSystematic ReviewNA MethodsQuery Search (141)Query Search (198)Title & Abstract Scan (84)Title & Abstract Scan (103)Full Text Scan (97)Top Cited (20)Wider SearchState-of-the-Art (Section 4.1)Experimental Evaluation (Section 4)
Figure 1: The search and filtering method with the number of remaining papers between brackets.
2.2 Results
Figure 1 summarizes the initial search results from WoS and Scopus, including duplicates. After filtering and removing duplicates, 97 papers remained, of which 11 were review papers. Classification based on Table 1 is detailed in Tables LABEL:tab:literature and Table LABEL:tab:literature2, with review papers excluded from category-specific analysis due to their broader scope.
Figure 2 illustrates the increasing trend in AML-related research, with additional category-specific trends shown in Appendix B. Table 2 lists journals and conferences with more than one relevant publication, covering 25 papers; the remaining 72 papers were distributed across other venues with a single publication each.
Figure 2: The number of papers incorporated in the full query and after filtering per year.
Crypto-related research accounts for 26 papers ($26.5\%$ of the total), analysed separately in Appendix C due to its distinct AML setting, data characteristics, and anonymity challenges. These differences are evident in the corresponding plots.
Appendix D presents results for the 20 top-cited papers, summarized in Table 4. The category distribution aligns with the overall literature, with $25\%$ focusing on cryptocurrency, closely matching the full-scope percentage.
Table 2: Journals/conferences that have two or more publications in scope of this review.
Journal/Conference Title Number
Journal of Money Laundering Control 6
International Conference on Advances in Social Network Analysis and Mining, ASONAM 3
CEUR Workshop Proceedings 2
IEEE International Conference on Data Mining Workshops 2
IEEE International Conference on Big Data (Big Data) 2
Journal of Physics: Conference Series 2
EPJ Data Science 2
Information Sciences 2
International Conference on Machine Learning Technologies 2
IEEE Access 2
3 Discussion on the Literature
3.1 Review Papers
We identified 11 review papers in scope for our literature review. These are analysed first to illustrate how our work complements and extends the existing literature.
Ngai et al. [100], Kurshan and Shen [74] and Lokanan [84] focus on financial fraud in general, with less emphasis on anti-money laundering or graph learning. Ngai et al. [100] construct a framework to classify financial fraud, but includes only one paper on money laundering, which was also the only paper on network analytics. Kurshan and Shen [74] give a high-level overview of the general difficulties of machine learning solutions for fighting financial crime. They noted that network-based methods have a lot of potential to process many transactions handled by financial institutions. The paper is closer to a discussion paper than to a systematic literature review. The scope of the review by Lokanan [84] is limited. It deals with both credit card fraud and money laundering, but only incorporates visualisation methods used for outlier detection.
Other review papers deal specifically with AML [75, 115, 42]. The work by Kute et al. [75] is most like ours. The authors give an overview of the methods applied in the literature, but their main topic is interpretability. Hence, their focus is less on the wider AML literature. Semenov et al. [115] discuss the different methods applied to AML. It touches upon different network methods, but it is not a systematic literature review. The framework by Gao and Ye [42] again deals with a wider array of methods, with the authors concluding that network analytics is the most promising. However, the number of papers on network analytics for AML included is limited.
Gerbrands et al. [47] look at the effects of money laundering policies, by measuring the change in network features when a major AML law is introduced. It analyses cluster size and the degree, closeness and betweenness centrality. Hence, the scope is different, since we focus on enhancing anti-money laundering using network features.
A significant number of review papers deal with crypto currency related money laundering [64, 27, 4, 56]. Irwin and Turner [64] describe challenges in detecting money laundering on the blockchain. As a result, the scope is smaller, with a total of 45 unique references. These different challenges are partially addressed by Alarab et al. [4] and Han et al. [56]. The experiment by Alarab et al. [4] compares different supervised machine learning methods, with the statistics of the ego-network of the different nodes in the Bitcoin transaction network [37, 139] as features.
Other papers cover a specific money laundering method and propose tailor-made solutions. Han et al. [56] discuss the application of anomaly detection, based on crime-specific transaction patterns. The paper also covers Ponzi schemes and blackmail campaigns. Their scope is more limited, with a total of 33 references. In the same line, Day [27] gives an overview of different characteristics of financial crime, and uses these to construct knowledge graphs for building legal cases.
We extend on prior work by conducting the first systematic review of network analytics methods for anti-money laundering, including both supervised and unsupervised methods. Our comprehensive scope spans traditional centrality measures to advanced graph neural networks, and addresses money laundering in both fiat and cryptocurrency contexts.
3.2 Full Scope of Papers
We start by analysis the 83 non-review papers. Although network analytics for anti-money laundering has been researched for a long time, especially after 2010 a lot of new research has been published. Figure 2 illustrates this increasing research interest, with an average increase in output of $61\%$ year-on-year for the papers in scope. Below, we dive deeper into the different aspects of these papers.
3.2.1 Methods.
The methods in the literature are supervised, unsupervised, semi-supervised and visualisations. The most popular method is unsupervised learning, followed by supervised learning and visualisations, as shown in Figure B1. The popularity of unsupervised learning may be due to money laundering being very uncommon, resulting in highly skewed data sets with very few labelled observations [3]. This is exacerbated by the limited resources institutions have to investigate the massive amounts of transactions, leaving most transactions not investigated [133, 107]. In addition, criminals try to evade detection by continuously changing tactics [10, 107], making historically suspicious patterns less relevant for future predictions [15]. The performance of (supervised) models trained on historical data can therefore decrease over time [14].
Diving deeper into the different modelling methods presented in Table LABEL:tab:literature2 and Figure B2, we see that half of the papers rely on manual feature engineering. These manual features often involve the classic network centrality measures (degree, betweenness etc.) and summary statistics on the transaction in the nodes’ ego-network. The second is clustering, which is broadly speaking used in two ways. On the one hand, clustering is applied for visualisation to avoid visual cluttering and give the user the possibility to select the level of detail shown. On the other, clusters are used to construct additional features to be used by detection models. Next to manual feature engineering and clustering, many methods can be seen as rule-based. All these insights show that the current literature still heavily relies on expert knowledge (to construct features and/or rules). Hence, there is a large opportunity to further explore and develop more data-driven machine learning methods in the fight against money laundering.
The evolution of methods used over time, as presented in Figure B2(b), shows that the earliest research relied almost exclusively on rule-based systems and manual feature engineering. These remain popular approaches to this day. From 2018 onward, we see the application of deep learning methods based on shallow representations and graph neural networks. Especially deep representations based on GNNs have seen a spike in 2022, the final year covered in this review.
The downstream methods used for classification have also evolved. Tree-based methods, clustering, anomaly detection and logistic regression have been used in earlier literature and are still in use today. However, we see that SVM-based methods were popular until 2019, while neural networks were only used in AML as of 2020. These are often recurrent neural networks to incorporate the temporal aspect of the networks [145, 154, 96, 81].
Key to these methods is also how the authors define the network(s) used. The construction of multiple networks, e.g., same entities but using different relations, to capture richer data is present in 18 papers. Of these papers, five used the undirected network or the one with direction reversed in tandem with the original network. It allows to propagate network information from, e.g., the receiver of a transaction back to the sender.
3.2.2 Objectives.
We identified multiple objectives of the methods for anti-money laundering. Figure B3 shows that the objective is most often detecting suspicious clients, followed by detecting suspicious transactions, money flows and communities. Client classification combines payments made over a longer period and the person’s characteristics. This seems better for detecting money laundering, since it is often done over a longer period, with the aim to make the individual transaction appear normal. Therefore, models trained on a series of transactions, leveraging an individual’s behaviour, can obtain higher performance.
Figure C3 shows that methods for crypto-currency are most often introduced for suspicious transaction detection. Due to the pseudo-anonymity, it is much harder to know which wallets belong to what person [157, 150], making it less feasible to build profiles. Therefore, researchers in AML have also looked at de-anonymisation of the wallets [157, 150, 48]. They use network features to cluster wallets together to see if they belong to the same person or organization.
In recent years, more studies have been done on flow/chain detection, tracking payments over multiple steps/people. Although this requires more computing power, it more closely resembles reality [130, 77]. Hence, methods leveraging these different steps and actors are becoming more widely used.
3.2.3 Evaluation.
The most popular metrics to evaluate the methods are the recall, $F1$-score, precision and time, as shown in Figure B4. We see that training/prediction time was a popular method in earlier studies, but relatively few papers use time in more recent work, which might be a consequence of the increase in computing power. Additionally, the popularity of the precision, recall and $F1$-score is a recent and increasing trend in the literature.
We remark two surprising results in our analysis. First, only one paper used the area under the precision-recall curve (AUC-PR). Due to the high label imbalanced, this metric is shown to be more appropriate than the area under the ROC curve (AUC-ROC) for evaluating performance [26, 103]. Therefore, we would expect it to be more widely used. Second, some papers solely reported accuracy [31, 146, 48, 148]. Given the class imbalance and the dependence on a threshold, accuracy fails in providing an appropriate performance assessment.
3.2.4 Data.
Figure B5 shows that most data sets are open-source, although proprietary data is also used in almost $50\%$ of the papers. Only a minority of papers report their findings using synthetic data. There is, however, a major difference between crypto- and other research. Looking at Figure C1, we see that crypto-research is almost exclusively done via open-source data. This results from the blockchain’s freely accessible distributed ledger [139]. On the other hand, classic transaction data cannot be shared, due to privacy reasons, resulting in half the data sets being proprietary.
Following this last remark, one would expect to have (almost) no research done on open-source data sets for fiat currencies. Figure C1 shows that almost $30\%$ of non-crypto related papers use open-source data. These papers can be categorised in four groups:
  • Methods that are evaluated on non-financial network data, whether or not next to proprietary/synthetic money laundering data: Zhou et al. [156], Ovelgönne et al. [102], Micale et al. [93], Prado-Romero and Gago-Alonso [109]
  • Methods evaluated on government data, assumed or stated to be open-source: Didimo et al. [29], Imanpour et al. [63], Velasco et al. [136], Bahulkar et al. [11], Malm and Bichler [88]
  • Methods constructed to analyse leaked documents/scandals, e.g., the Panama papers: Joaristi et al. [69, 70], Magalingam et al. [85], Adriaens et al. [1], Winiecki et al. [141], Helmy et al. [59], Cheong and Si [23]
  • Other: Nandhini and Das [98], Bhalerao et al. [13], Çağlayan and Bahti̇yar [17]
For the Other category, we have the following. Nandhini and Das [98] say that the data is publicly available, without giving any more information. Bhalerao et al. [13] scraped their data from Hack Forums. Finally, Çağlayan and Bahti̇yar [17] use a Kaggle data set that is based on synthetic data, which we categorise as open-source, since it is made publicly available.
We present the open-source data sets in Table 3. The most widely used are almost all crypto-currency data sets. Raw Bitcoin data was used by ten papers. This raw data is often enhanced using laundering data from other platforms. One such popular platform is WalletExplorer, which is used by three papers. It is less relevant today, since the data has not been maintained since 2016. Another recent data set on Bitcoin, which is becoming more widely used, is the Elliptic data sets [37, 139]. Although only published in 2019, it has already been used in six experiments. After Bitcoin, Ethereum seems to be the second most popular crytpo-currency for AML practices.
The most used data set for fiat currency is the Czech Financial Data set (CFD). This data set does not provide any labels. The papers using this data set introduce additional, synthetic money laundering patterns to test the methods [80, 138]. This is not ideal, as this hinders comparison across research and can introduce detection bias.
As mentioned above, AML methods are also being evaluated on leaked data like the ICIJ Offshore Leaks Database (including, e.g., the Panama-papers) and the Enron data set. However, these are less interesting when evaluating the performance on financial data.
Table 3: The open-source data sets used in the different papers.
Data Set Number References
Raw bitcon 10 [89], [142], [106], [143], [129], [112], [150], [125], [111], [151]
Elliptic Data set 6 [5], [145], [96], [122], [78], [81]
Ethereum XBlock.pro 4 [125], [157], [66], [67]
ICIJ Offshore Leaks Database 4 [70], [44], [69], [141]
WalletExplorer 3 [142], [129], [150]
Czech Financial Data set (CFD) 2 [80], [138]
Enron Data set 2 [85], [1]
RenRenDai 1 [146]
Bloxy 1 [67]
iCOV 1 [47]
Secret Presidential Funds scandal 1 [23]
Tencent QQ 1 [159]
Caviar data set 1 [11]
BitcoinTalk 1 [143]
Ripple Network 1 [19]
2015 Ashley Madison extortion scam 1 [106]
FinCEN 1 [29]
BlockCypher 1 [111]
3.3 Top Cited Papers
A final step in our discussion is an in-depth analysis of the top-cited papers. These papers are well received by the scientific community and are likely of high quality. We select the $10\%$ most cited papers per year, extended with the $10\%$ most cited papers overall (see Section 2.1). This results in 20 papers as listed in Table 4. The analysis is done in two stages to gain deeper insights. First, we revisit the different categories as presented in Table 1. Second, the papers are classified according to the processing steps and challenges, as introduced by Bockel-Rickermann et al. [14].
Table 4: Summary of the top-cited papers, with indication of whether it is a review and/or crypto-related.
Reference Title
Year
Citations
Review
Crypto
Ngai et al. [100] The application of data mining techniques in financial fraud detection: A classification framework and an academic review of literature 2011 1307
Fronzetti Colladon and Remondi [41] Using social network analysis to prevent money laundering 2017 200
Chang et al. [22] WireVis: Visualization of Categorical, Time-Varying Data From Financial Transactions 2007 171
Senator et al. [116] Financial Crimes Enforcement Network AI System (FAIS) Identifying Potential Money Laundering from Reports of Large Cash Transactions 1995 152
Dreżewski et al. [33] The application of social network analysis algorithms in a system supporting money laundering detection 2015 141
McGinn et al. [89] Visualizing Dynamic Bitcoin Transaction Patterns 2016 122
Gao and Ye [42] A framework for data mining-based anti-money laundering research 2007 106
Wu et al. [142] Detecting Mixing Services via Mining Bitcoin Transaction Network With Hybrid Motifs 2021 94
Malm and Bichler [88] Using friends for money: the positional importance of money-launderers in organized crime 2013 66
Zhou et al. [156] A Local Algorithm for Structure-Preserving Graph Cut 2017 63
Li et al. [80] FlowScope: Spotting Money Laundering Based on Graphs 2020 54
Irwin and Turner [64] Illicit Bitcoin transactions: challenges in getting to the who, what, when and where 2018 47
Phetsouvanh et al. [106] EGRET: Extortion Graph Exploration Techniques in the Bitcoin Network 2018 38
Kute et al. [75] Deep Learning and Explainable Artificial Intelligence Techniques Applied for Detecting Money Laundering–A Critical Review 2021 38
Zhdanova et al. [155] No Smurfs: Revealing Fraud Chains in Mobile Money Transfers 2014 32
Bhalerao et al. [13] Mapping the underground: Supervised discovery of cybercrime supply chains 2019 24
Ovelgönne et al. [102] Covertness Centrality in Networks 2012 20
Cheong and Si [23] Event-based approach to money laundering data analysis and visualization 2010 12
Gerbrands et al. [47] The effect of anti-money laundering policies: an empirical network analysis 2022 11
Zhou et al. [157] Behavior-Aware Account De-Anonymization on Ethereum Interaction Graph 2022 7
3.3.1 General Classification.
The results show that five out of 11 review papers are selected. This is to be expected, since review papers have a higher likelihood of being cited [32]. Plots summarising the classification according to Table 1 are given in Appendix D. Similar results are observed for the top-cited papers as for the literature as a whole; unsupervised learning is by far the most applied method, and the research relies heavily on manual and rule-based features. For the evaluation of the methods, only time and threshold-dependent metrics were used. This indicates that threshold-independent metrics are not fully mainstream when comparing AML models using networks analytics. We observe that, compared to the literature as a whole, flow/chain detection methods are better cited as well as studies using open-source data. As mentioned before, flow detection better captures the intricate patterns used by money launderers. The use of open-source data makes comparison with and replication of the studies possible, fostering adoption of the presenting research.
Additionally, five papers deal with crypto-related money laundering, indicating a strong research interest. The literature mentions two causes explaining the growing interest. On the one hand, crypto currencies are widely used by criminals because of the absence of regulation [40] and the anonymity it provides [64, 139]. On the other hand, a full history of ownership is available on the blockchain for each coin, facilitating the discovery of transaction patterns in payment networks [64, 139].
Since this paper discusses network learning, an additional summary of the construction of the networks is given in Table 5. Most of the papers use homogeneous networks where transactions form the edges between accounts. Only a few use heterogeneous networks, either by having transactions as nodes as well, or having nodes that represent different characteristics or feature values.
Table 5: Summary of the construction of the network.
Paper Nodes Edges
Fronzetti Colladon and Remondi [41] Factoring companies Economic sector geographical data transaction amount tacit link (same owners/resources)
Chang et al. [22] People Transactions
Senator et al. [116] Accounts People, businesses, accounts and location Connections
Dreżewski et al. [33] People (via court register) Transactions (via bank statements)
McGinn et al. [89] Transactions, inputs and outputs Input/output part of same transaction, inputs that belong to the same address
Wu et al. [142] Address Address and transactions Transactions If they are related
Malm and Bichler [88] People (from police reports) Co-appearance in the report (family, client-lawyer etc.)
Zhou et al. [156] All items of interest (bank account, name, email, address, phone number) Link bank account to other items
Li et al. [80] Bank accounts Transfers
Phetsouvanh et al. [106] Bitcoin wallet address Transactions
Ovelgönne et al. [102] - -
Cheong and Si [23] People Relations
Zhou et al. [157] Accounts Transactions
Zhdanova et al. [155] mWallets Transactions
3.3.2 Processing steps.
Next, these papers are classified according to the framework presented by Bockel-Rickermann et al. [14]. Table 6 summarises the papers over the different processing steps. The most prevalent pre-processing steps are feature engineering and exploration. The feature engineering is done to define different risk profiles or to enhance GNN models. We see in the literature that this is either based on neighbourhood and centrality metrics [41, 33, 102] or on the aggregation of local (node-specific) information [157, 155]. For exploration, some of the research is done on how to construct meaningful visualisations [22, 89], while other papers rely on clustering the nodes as additional support for their findings [88, 41].
Most papers apply unsupervised learning in their processing step. This often involves detection of suspicious flows of transactions [156, 80, 106]. Another important part of the literature uses centrality metrics to define specific roles for entities in the network [33, 88, 102]. Finally, unsupervised learning methods can involve creating heuristics to define outlying behaviour [116, 22, 23, 155].
For the post-processing step, most research covers statistical evaluation, implementation and interpretation. The statistical evaluation is used to have a quantitative evaluation of the performance of the method. All research on the implementation looks at the efficiency of the methods. It deals solely with computation time and scalability. To have interpretations in the post-processing step, researchers use specific interpretable features.
Table 6: Papers per method cluster, based on Bockel-Rickermann et al. [14].
Process Step Sub-Steps References
Pre-processing Sampling [41].
Exploration [41],[22],[89], [88].
Missing Value Treatment [41].
Outlier Detection and Treatment -
Categorization, standardization & segmentation [33], [13].
Feature Engineering [41], [33], [102],[157], [155].
Variable Selection -
Processing Unsupervised Learning [22], [116], [33], [88], [156], [80], [106], [102], [23],[155].
Supervised Learning [41], [13].
Semi-Supervised Learning [142].
Hybrid Learning [157].
Post-processing Statistical Evaluation [41], [142], [156], [80], [13], [157], [155].
Interpretation [33], [142], [13], [157].
Economical Evaluation [155].
Implementation [33], [89], [156], [80], [102],[157], [155].
More detailed trends in the literature are discussed below according to three main categories, that mainly follow the processing steps; unsupervised learning, (semi-)supervised learning and visualisation methods.
Unsupervised learning can be divided into two main streams. The first stream assigns and identifies specific roles of actors in criminal networks, based on police reports covering past criminal cases [88, 33]. The second stream introduces methods to track money flows through the network [156, 80, 106]. It is important to integrate these transaction chains, since money laundering is a process involving multiple steps and actors [77]. Next to these two main streams, the remaining paper [102] introduces a new covertness measure.
The above-mentioned methods often rely on the application of centrality metrics, e.g. degree and betweenness centrality, which are combined and interpreted in different ways, leading to tailor-made solutions for specific sub-problems of anti-money laundering methods. Although detecting money laundering is in principle trying to find anomalous behaviour, very few studies deal with (unsupervised) anomaly detection [86, 109, 34], leaving many possibilities for future work.
(Semi-)Supervised learning methods deal with different aspects of money laundering. Their strength often lies in a combination of (1) the methods used and (2) tailoring the network to the problem at hand. The obtained network features are put into downstream classifier. To maintain interpretability, network metrics are recalculated for multiple networks, each representing a specific aspect of the data, to extract predictive features [41, 142]. Promising methods are often based on deep learning [157], but these are mostly black boxes. However, research on the interpretability of deep network representation methods, as mentioned by Kute et al. [75], is still scarce.
Visualisation of transaction networks also gets a strong focus in the literature [116, 29, 22, 23, 89]. The general idea is to support the experts in performing their investigations. This is either done by having a overview of the flow of money over different transactions [89, 22] or by finding specific relations among the persons involved [116, 29, 23]. Hence, these methods are mostly intended to be used after suspicion has been raised. None of them include evaluation metrics to test their strength as money laundering detection methods.
3.3.3 Challenges.
Table 7 summarises the papers according to the challenges identified by Bockel-Rickermann et al. [14]. It only contains those challenges that were mentioned in the literature. The main challenge is data availability, followed by unlabelled data. Due to privacy issues, only a limited number of anonymised data is available. Additionally, since transaction data sets are often huge and resources are limited, it is hard to provide all instances with a label. This is the main obstacle for the construction of a general benchmark. On the other hand, few papers mention feature construction, real-time execution, or verification latency as a challenge. The latter two are important when detecting transaction fraud, where decisions need to be made within seconds [133, 131]. For anti-money laundering, however, the investigation is always done after the fact, based on a longer history of transactions [77]. Hence, it is less reliant on execution time.
We introduce four additional challenges observed in the anti-money laundering literature, supplementing the work of Bockel-Rickermann et al. [14]. We view this as an extension of the framework since these challenges can be important to fraud research in general.
  • Bias in data affects the performance of the model and manifests itself in different ways. First, there is bias in the missing values [88]. Second, when using synthetic data, there could be detection bias [155] present, when the synthetic data includes specific patterns that the model is trained to look for.
  • Generalisation is mainly used in two ways. One way refers to the ability for methods specifically developed for anti-money laundering to be used for other fraud domains as well [157]. Another and important use of the term refers to the need for methods to generalise to unseen data [49], often referred to as inductive methods (compared to transductive ones). This is an important consideration when selecting network-based methods [55, 131, 132].
  • Robustness of a method is important since money laundering can be seen as an adversarial attack [80], as perpetrators try to avoid detection by the methods.
  • Anonymity of the account data gives rise to two main challenges. First, the identity of the account holder is unknown in crypto data sets, which makes it almost impossible to know if different accounts - also called wallets - belong to the same person [142]. Second, financial institution possess customer data, but due to privacy reasons, researchers cannot publish these. Hence, published articles avoid using Personally Identifiable Information all together [13].
Table 7: Papers per fraud detection challenge, based on Bockel-Rickermann et al. [14].
Challenge Papers
Original Automation [116], [23], [157].
Class Imbalance [142], [13].
Concept Drift [41], [13].
Data Availability [41], [22], [142], [88], [102], [155].
Feature Construction [102].
Noisy Data [116], [142], [88].
Stream Data [116], [89], [157].
Real-Time Execution [41].
Scalability [156], [102], [157].
Unlabelled Data [116], [142], [13], [157].
Verification Latency [155].
New Bias in data [88], [155].
Generalisation [157].
Robustness [80].
Anonymity [89], [142], [13].
3.4 Further Considerations
Based on the observations above, we give an intermediate conclusion and perspective on future work. The methods in the literature as a whole still rely heavily on expert-based methods, i.e., manual feature engineering and rule-based systems, while deep learning methods both for network representation as for down-stream classifiers are gaining traction. A more detailed analysis of these methods is needed to evaluate their performance properly, especially given that the current literature barely applies threshold-independent metrics meant for highly skewed data, like the AUC-PR. Moreover, our analysis showed that none of the top-cited papers apply threshold-independent metrics. This detailed analysis is what we are going to do in the second part of this paper.
Different gaps in the literature still exist which we believe gives ample opportunities for future work. A first gap was also identified by Kute et al. [75], namely the lack of interpretation tools. Future work can focus on the different interpretation methods that already exist to try and explain the output of graph neural networks.
A second gap comes from the fact that classic unsupervised methods [21, 3] do not seem to be well represented. Future work should analyse the effectiveness of unsupervised learning in networks to find anomalous behaviour. This can be beneficial since many clients and transactions are effectively unlabelled, and the modi operandi for money laundering keeps evolving. The second part of this paper will introduce a first way of applying such unsupervised learning methods, but a more extensive study is needed.
A third is that most methods are based on the analysis of static networks. Some research is being done to incorporate the time-dynamic nature of transaction networks using recurrent neural networks, but we believe that much more can still be done here.
A fourth gap is related to the inherent incompleteness of labels in AML. As criminals try to cover their tracks and financial institutions have only limited resources [133, 107], it is a given that there are still many money laundering cases that stay unnoticed. These transactions are seen as having label 0 when training the models, resembling more a positive and unlabelled (PU) problem. Specific PU-learning methods for networks should be applied when training a model.
A fifth gap is related to the limited resources as well. Since not all clients and transactions can be investigated, the predictions should take this into account. One way of dealing with this is learning-to-rank [28, 135], which has already been shown to be effective in fraud detection. The main benefit is that we focus more on correctly ranking the top most suspicious clients, while less importance is given to have a correct ranking for less suspicious clients. Although this is highly relevant, it has not yet been applied to the AML literature using network analytics.
4 Experimental Set-Up
Most papers discussed above use centrality measures, while only few incorporate deep learning. To the best of our knowledge, a benchmark study comparing the state-of-the-art network learning methods for anti-money laundering is still missing. This section describes the experimental framework to fill this gap, including model specification (Section 4.1), data (Section 4.2), hyperparameter tuning (Section 4.3) and performance metrics (Section 4.4). The focus is on supervised learning, given that the data sets used contain ground-truth labels. While some unsupervised methods are explored, a comprehensive experiment is left for future work. Visualization methods are excluded as they serve investigative support rather than objective classification. Given challenges in open-source data availability, we present results on two datasets and provide code to facilitate reproducible experiments. We hope this framework will support and expand future research, and allow uniform publishing and comparing of results on proprietary data without the need to disclose sensitive information.
4.1 Model Specification
To set up the experimental evaluation, we start from the wider literature on network representation. Extensive overviews are given by Cai et al. [18], Hamilton et al. [55] and Goyal and Ferrara [50], classifying the methods according to their input, output, and embedding technique. A sub-branch of representation learning applies deep learning methods [158, 144]. Based on these papers, we select the methods that are most prominently used. We refined the selection by considering that, for AML, transaction data contains networks with millions of nodes and edges, making scalability an important selection criterion. Table 8 provides an overview of the methods selected. They are grouped into three categories [18, 55, 50]: manual feature engineering, shallow representation learning and deep representation learning. These methods will be tested against a baseline model that only includes the intrinsic features, denoted by IF.
Table 8: The methods used to generate network features for the experimental evaluation.
Category Features Definition
Manual feature engineering Density Baesens et al. [10]: Number of edges in the egonet relative to the maximal number possible.
Centrality measures Closeness, betweenness and eigenvector centrality.
PageRank Page et al. [104]: Score indicating the importance of a node.
Shallow representation learning DeepWalk Perozzi et al. [105]: Random walk through the network. Embedding via word2vec.
Node2vec Grover and Leskovec [51]: Truncated random walks (breadth-first vs. depth-first). Embedding via word2vec.
Deep representation learning GCN Kipf and Welling [71]: Graph convolutional network
GraphSAGE Hamilton et al. [55]: Aggregation based on fixed-size sample of neighbours
GAT Veličković et al. [137]: Graph attention network
GIN Xu et al. [149]: Graph isomorphism network
Manual feature engineering includes both local and global metrics. The first is the density of the ego-network. A node’s ego-network consists of said node, its direct neighbours and all connections between the neighbours. The density is the relative number of connections in this ego-network compared to the theoretical maximum. Summary statistics, i.e., minimum, mean and maximum values, of the density of the node’s neighbours are also included.
A second type of manual features are centrality metrics, quantifying the global position of the node in the network [99]. The most popular centrality metrics are the betweenness, closeness and eigenvector centrality. The betweenness quantifies the number of shortest paths between any two nodes in the network passing through a given node. The closeness measures how close a node is to all other nodes, and is the inverse of the average distance to the other nodes. The eigenvector centrality is high for nodes that are connected to (but are not necessarily themselves) important nodes in the network. The importance of the node itself can be quantified using the PageRank [104].
Next to the manual feature engineering, automatic network features are constructed using network embedding methods [55, 50, 18]. Although there are many different methods to construct a network embedding [18], we focus on those based on deep learning, often called network representation learning [131, 133]. Network representation learning comprise the second and third category of features in Table 8.
Shallow network representation learning can be viewed as an “embedding lookup” [55, 133]. Our experimental set-up will use different methods based on random walks. DeepWalk [105] was the first method to try to give nodes that are close in the network, i.e., that co-occur on short random walks, a similar embedding. A random walk is seen as a sentence of words, which can be embedded into a Euclidean latent space using NLP methods, often via skip-gram [95].
Node2vec [51] extends the capabilities of the random walks by introducing two additional hyperparameters, $p$ and $q$, resulting in the option to either use breadth-first or depth-first sampling. The method samples each node according to the following unnormalized transition probability, $\alpha$, at node $v$, when coming from node $t$:
$\alpha(c_{i+1}=x\mid c_{i}=v,c_{i-1}=t)=\begin{cases}\frac{1}{p}&\text{if }d_{tx}=0\\ 1&\text{if }d_{tx}=1\\ \frac{1}{q}&\text{if }d_{tx}=2\\ \end{cases},$
where $d_{tx}$ denotes the distance between node $t$ and $x$. These probabilities are only defined when there is a link from $v$ to $x$. Otherwise, it is set to 0. Depending on the values of $p$ and $q$, the embedding focuses more on capturing either homophily or structural equivalence [51].
The main drawbacks of shallow representation learning methods, are that (1) they are transductive, so need to be retrained fully when a new node is added to the network, and (2) retraining on the same network often results in a completely different embedding. This second point can be understood intuitively by viewing two embeddings of the same network, but where one is a rotation of the other. Both are equally good, but a downstream classifier trained on one cannot be used to make predictions on the other, as the coordinates, i.e., the features values, are completely different.
This needs to be considered when training and comparing models. In the experiments presented below, we construct the embedding first on the network with nodes in the training and validation set, only training on the labels from the training set. Afterwards, the embedding is constructed for the full network, and we train a new classifier using the labels of the train and validation set.
This mimics the applications of these methods in reality, since a bank will have the full transaction network up to the current time, with some historical labels from past moments. Based on this, they need to determine which accounts to investigate next for money laundering. This also highlights the difference with credit card fraud, where a new transaction comes in and a decision needs to be made in only a few seconds, limiting the ability to retrain the embedding on the whole network.
Deep network representation learning adapts deep learning methods to be directly applicable to networks. This paper applies different graph neural network (GNN) architectures. The first graph neural network framework was introduced by Scarselli et al. [113] and Scarselli et al. [114] and is based on recurrent neural networks. The idea is that the network defines how nodes can propagate information to each other. This propagation step is iterated until a state of convergence. As noted by Hamilton et al. [55], the initial message passing idea is mostly used for sub-network embedding calculation.
The most popular methods for node embedding calculation are graph convolutional networks (GCNs), based on convolutional neural networks [71]. It relies on neighbourhood aggregation to update the embedding as follows:
$H^{(l+1)}=\sigma\left(\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}H^{(l)}W^{(l)}\right),$
with $H^{(l)}$ the embedding at step $l$, $\sigma$ a non-linear (activation) function, $\tilde{A}=A+I$ the adjacency matrix with self-loops added, the diagonal matrix $\tilde{D}_{ii}=\sum_{j}\tilde{A}_{ij}$, and $W^{(l)}$ the learnable weights in layer $l$. This means that the node embedding for the individual node is updated via
$h_{i}^{(l+1)}=\sigma\left(\sum_{j\in\tilde{\mathcal{N}}_{i}}\frac{1}{\sqrt{\text{deg}(i)\text{deg}(j)}}W^{(l)}h_{j}^{(l)}\right).$
Neighbour aggregation has seen multiple extensions, two of which this paper will evaluate. The first is GraphSAGE [55], which samples a fixed number of neighbours for each node, to make the calculations scalable to large graphs. The representations of the sampled neighbours are aggregated, with the authors proposing the mean, LSTM and pooling aggregators [55]. Then, the aggregated neighbourhood representation is concatenated with the node’s representation. Finally, the representation is updated using the weights and activation function. For a given node $v$, this is calculated as:
$h_{\mathcal{N}_{v}}^{(l)}$ $=$ $\texttt{AGGREGATE}_{l}\left(\{h_{u}^{(l-1)},\forall u\in\mathcal{N}_{v}\}\right)$
$h_{v}^{(l)}$ $=$ $\sigma\left(W^{(l)}\cdot\left[h_{v}^{(l-1)}\mathbin{\|}h_{\mathcal{N}_{v}}^{(l)}\right]\right),$
[…]
The second extension is graph attention network (GAT) [137], which adds an attention mechanism when learning the representations. Our work uses the modified version of the attention mechanism, introduced to fix the static attention problem [16]:
$e_{ij}$ $=$ $a^{T}\texttt{LeakyReLU}\left(W^{(l)}\left[h_{v}^{(l-1)}\mathbin{\|}h_{\mathcal{N}_{v}}^{(l)}\right]\right)$
$\alpha_{ij}$ $=$ $\texttt{softmax}_{j}(e_{ij})$
$h_{i}^{(l)}$ $=$ $\sigma\left(\sum_{j\in\tilde{\mathcal{N}}_{i}}\alpha_{ij}W^{(l)}h_{j}^{(l-1)}.\right)$
The above formulas apply one attention mechanism. As mentioned by Veličković et al. [137], having multiple mechanisms - also called heads - can improve stability. We will follow the method of Veličković et al. [137], where the intermediate steps concatenate the results over the heads, while the final layer averages the results.
A final graph neural network method selected in this paper is the Graph Isomorphism Network (GIN) [149]. The authors used the Weisfeiler-Lehman graph isomorphism test to define the necessary conditions a GNN must satisfy to achieve maximal discriminative power. This discriminative power is denoted by how well the training data is fitted.
Using the universal approximation theorem [61, 60], GIN updates the representation using a multilayer perceptron (MLP).
$h_{v}^{(l)}=\texttt{MLP}^{(l)}\left(\left(1+\epsilon^{(l)}\right)h_{v}^{(l-1)}+\sum_{j\in\mathcal{N}_{i}}h_{j}^{(l-1)}\right)$
Although this increases the risk of overfitting, the authors show that GIN has satisfactory generalisation capabilities. The main problem with other GNNs is that they tend to underfit the data [149]. We include GIN because recent research showed its potential to outperform other GNN architectures [36, 127].
There are some key differences between the three categories presented in Table 8. As mentioned before, the manual features need to be specifically defined, while representation learning tries to find meaningful embeddings automatically. Another difference concerns the node-specific or intrinsic features. The manual features and those from shallow representations are added to the intrinsic features, which are used in a down-stream classifier (as illustrated in Figure 3). The GNNs, on the other hand, directly incorporate the intrinsic features. As a result, the embeddings are directly learned using the intrinsic features and the classification task. Hence, shallow representation learning is closer to unsupervised learning, only based on the network topology, while deep representation learning are supervised learning methods.
ID F1 F2 $\ldots$ Fn
A f11 f21 $\ldots$ fn1
B f12 f22 $\ldots$ fn2
$\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$
E f1k f2 $\ldots$ fnm
ID F1 F2 $\ldots$ Fn NF1 $\ldots$ NFm
A f11 f21 $\ldots$ fn1 nf11 $\ldots$ nfm1
B f12 f22 $\ldots$ fn2 nf12 $\ldots$ nfm2
$\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$
E f1k f2 $\ldots$ fnm nf1k $\ldots$ nfmk
ABCDENetwork Embedding
Figure 3: For the feature engineering and shallow representation methods, the obtained network features/embedding is added to the intrinsic ones.
Different decoders are used for the final classification task in the supervised setting. A deep neural network with two hidden layers, each of dimension ten is used for the manual features and shallow representation methods. For the graph neural networks, the GNN layers are followed by a single linear layer for making predictions.
For the unsupervised model, we apply isolation forest [82], as it achieves state-of-the-art performance for outlier/anomaly detection [126, 65]. This will be applied to the intrinsic and manual features, and the shallow representations, since the isolation forest requires a tabular input. As mentioned before, a larger experiment comparing unsupervised methods for network analytics is left for future work.
One of the objectives of this research is to provide the code for all methods in a clean and simple way and make these methods easily accessible to researchers and practitioners. The methods are implemented in Python. The manual features are constructed using NetworkX [53] and NetworKit [7], and the representation learning is implemented in PyTorch Geometric [39]. The code is made available on GitHub [^1] https://github.com/B-Deprez/AML_Network.
4.2 Data
For the evaluation of the different methods, we rely on two data sets; the elliptic [37, 139] and IBM money laundering [6, 36] data sets. Using these two, we cover the two main lines of research in the literature. The elliptic data set concerns real-world crypto-transactions, while the IBM data set deals with simulated transactions in fiat currencies. We also discuss the main characteristics of and differences between the two networks based on different metrics.
4.2.1 Elliptic.
The Elliptic data set [37, 139] contains Bitcoin transactions for 49 time steps each being around two weeks long. We make this choice since (1) we know from Section 3.2 that this data set is widely used and (2) the instances have a ground truth label, facilitating evaluation of the methods. The network has 203 769 nodes and 234 355 edges. The nodes correspond to transactions, and an edge represents that the output of one transaction is the input of the next. Additionally, the data set contains 166 pre-calculated features. These are split into 94 transaction-specific, or local, features and 72 aggregated features, summarizing the local features of a node’s neighbours.
The Elliptic data set has only numerical features, which have already been standardised. The local features will be used for all methods. However, since the deep representation learning methods aggregate the neighbourhood information, including the node features, we do not include the aggregated features for these. This was also addressed by Weber et al. [139], stating that graph neural networks are better to address the heterogeneity of the neighbourhoods than the aggregated features.
Around $33\%$ of transactions are labelled, of which 4 545 are illicit and 42 019 are licit. Hence, the classification problem is imbalanced with around $11\%$ of nodes labelled, or $2\%$ of all nodes labelled illicit. We note that an illicit label does not automatically mean that the transaction is used to launder money. However, criminals will try to obscure the source of these illicit funds, so we believe that it is still informative.
We use the copy of the Elliptic data set included in the Pytorch Geometric library [25, 39]. We split the periods into a train (period 1-30), validation (period 31-40) and test set (period 41-49). Only nodes having a label are used for performance evaluation. Transactions with label 2 (unknown) are included for the construction of the network, but they do not contribute to training and loss calculation [71, 152, 139].
4.2.2 IBM Money Laundering.
The second data set is the HI-Small data set, taken from Kaggle, consisting of simulated transactions in a virtual world between individuals, companies and banks [6, 36]. A fraction of these transactions are labelled as money laundering. Given that the labels are at transaction level, the network is constructed with the transactions as nodes. Edges are added from transaction $v_{i}$ to $v_{j}$ if the receiver of transaction $v_{i}$ is the sender of transaction $v_{j}$, transaction $v_{j}$ happened before $v_{i}$, and the difference in time is smaller than $\Delta t$, i.e., $\Delta t\geq t_{j}-t_{i}\geq 0$. This is inspired by the work by Tariq and Hassani [124], and allows to capture the money flow. We take a 60-20-20 train-validation-test split based on transaction time.
Since this results in a massive data set and tuning all hyperparameter for DeepWalk and node2vec would take too long, we only take the last 500 000 transactions, and set $\Delta t$ to four hours. This results in a network with 500 000 nodes and 1 278 952 edges. Just as with the elliptic data set, this subsample of the IBM money laundering data set is highly imbalanced. Just 1356 transactions, i.e., $0.27\%$, are labelled as money laundering.
Each transaction has a couple of features, of which we use the following: ‘Amount Received’, ‘Receiving Currency’, ‘Amount Paid’, ‘Payment Currency’, ‘Payment Format’. Next to that, we extract the day of the week, hour and minute out of a transaction’s date stamp. The categorical features, i.e., ‘Receiving Currency’, ‘Payment Currency’, ‘Payment Format’, are transformed to numerical using one-hot-encoding. In the remainder of this paper, we denote this data set as IBM-AML.
An analysis of the topology of both networks is given in Table 9, based on well-known network features. IBM-AML seems to be much more connected, since the average degree is higher, and the average path length is lower compared to the Elliptic data set. This is corroborated by the degree distribution and the distributions of the closeness centrality, in Figure 4 and Figure 5, respectively. We see that the Elliptic data set more closely resembles a scale-free distribution, while the IBM-AML data set deviates from this with fewer nodes with low degree and more nodes with higher degree (hubs). The existence of hubs resulted in a much lower average path length, and higher closeness centrality values.
Table 9: Network features for the Elliptic and IBM-AML data set.
Network Avg. Degree Avg. Path Length Clustering
Elliptic 2.3 15.96 0.0117
IBM-AML 5.1 3.78 0.0000
Figure 4: Comparison of the degree distributions of the Elliptic and IBM-AML data sets.
Figure 5: Comparison of the distributions of the closeness centrality of the Elliptic and IBM-AML data sets.
4.3 Hyperparameter Tuning
Most methods are highly dependent on their hyperparameters. Hence, finding those that give the best performance is crucial. However, it is not feasible to tune these using a grid search. As the methods often have many hyperparameters, the number of combinations increases exponentially. Additionally, transaction networks are very large, which results in a relative long training time for each individual combination of hyperparameters. Therefore, we tune the hyperparameters efficiently via Optuna [2], whereby we give a range of possible values to select from. We use optuna’s default TPESampler sampler.
After these trials, the combination is retained that maximises the AUC-PR on the validation set. The ranges are summarised in Table 10 and 11, for supervised and unsupervised learning, respectively. These are based on the ones often used in the literature. As the embeddings from the shallow representation methods are used in both a supervised and unsupervised setting, the hyperparameters will be tuned for the two applications separately. This will be indicated with superscript S (for supervised) and U (for unsupervised) when presenting the results. A model only based on the intrinsic features (IF) is included. The number of trials is set to 50 for DeepWalk and node2vec, and 100 for the GNNs.
Since the IBM data set is much larger, scalability in both time and memory are becoming very important. Especially DeepWalk and node2vec take quite some time to run. The higher average degree also makes transition probability calculations more expensive. That is why the number of epochs for calculating the embedding are capped at 100. The attention mechanism of GAT is conditional on the query node, meaning that attention is learned for each node pair. To be able to train the model in memory, we limit the attention head to a maximum of two.
Table 10: The overview of the hyperparameter values for the supervised learning models. The tuned hyperparameter value is added between parentheses next to the name of the model.
Hyperparameter Tuning Range Model (hyperparameter value Elliptic $\mid$ IBM )
[…]
Latent/Embedding dimension $[32,128]\cap\mathbb{N}$ GCN ($108\mid 97$), GraphSAGE ($48\mid 53$), GAT ($99\mid 51$), GIN ($74\mid 47$)
[…]
GNN layers $[1,3]\cap\mathbb{N}$ GCN ($1\mid 2$), GraphSAGE ($1\mid 2$), GAT ($1\mid 1$), GIN ($1\mid 1$)
Number of neighbourhood samples Not included at the moment. $[2,5]\cap\mathbb{N}$ GraphSAGE (2)
[…]
Number of attention heads $[1,5]\cap\mathbb{N}\mid[1,2]\cap\mathbb{N}$ GAT ($4\mid 2$)
Dropout rate $[0,0.5]$ GCN ($0.288\mid 0.163$), GraphSAGE ($0.350\mid 0.244$), GAT ($0.186\mid 0.188$), GIN ($0.271\mid 0.273$)
[…]
Table 11: The overview of the hyperparameter values for the unsupervised learning models. The tuned hyperparameter value is added between parentheses next to the name of the model.
Hyperparameter Tuning Range Model (hyperparameter value Elliptic $\mid$ IBM )
[…]
The decoders used for classification all have a fixed architecture (Section 4.1). For the GNNs, the training of the decoder weights happens simultaneously as the other weights during training. For the manual and shallow representation features, the decoder/classifier needs to be trained afterwards. The only hyperparameter tuned is the number of epochs. The results are presented in parentheses in Table 10 and Table 11.
4.4 Performance Metrics
We tackle money laundering as a binary classification problem. We adopt the most popular performance metrics from the literature, namely the precision, recall and F1-score (Section 3.2). A threshold is needed for the output of the model to classify an instance as a money launderer. Since investigation resources are limited [133], the thresholds will typically be set relatively high, e.g., by classifying the top $0.1\%$, $1\%$ or $10\%$ of scores as money laundering. This is supplemented by a threshold equal to the relative prevalence of money laundering in the data sets.
The selection of these thresholds is ad-hoc, and can influence the conclusions. Therefore, we also include threshold-independent metrics. The most popular one for binary classification is the area under the ROC curve (AUC-ROC). The ROC curve plots the true positive rate against the false positive rate. However, since the distribution of labels is highly imbalanced, the false positive rate can be insensitive to changes in the number of false positive examples [26]. The area under the precision-recall curve (AUC-PR) is better suited to compare these models [26, 103]. This also adds to the literature, since we showed in Section 3.2 that few papers base their conclusions on the AUC-PR.
The performance metrics are calculated 100 times on a randomised version of the test set to get a grip on the variance of the results. The randomisation of the test set is done as follows. The manual network features and shallow representation embeddings result in classic tabular features (cf. Figure 3). For these, we apply a bootstrap method, where a new test set is constructed of the same size as the original, but where the observations are sampled from the original test set with replacement. The deep representation methods apply test masks. Therefore, we initialise new test masks by randomly selecting half the test masks from the original ones. Both ways of working result in 100 different test sets on which the performance metrics are calculated.
5 Results and Discussion
5.1 Supervised Learning
The results for the threshold-independent metrics, i.e., AUC-ROC and AUC-PR, on the test set are shown in Table 12, including the standard deviation calculated as described in Section 4.4. Additionally, the threshold-dependent results, i.e., precision, recall and F1-score, are given Table 13 and Table 14 for the top-$1\%$ results. To analyse the sensitivity to the threshold, the metrics are calculated when classifying the top $0.1\%$, $10\%$ and $p\%$ of scores as money laundering, with $p\%$ indicating the relative prevalence of the money laundering label in the training data set. The results for these three thresholds are included in Appendix E.
Table 12: Threshold-independent metrics: AUC-ROC and AUC-PR values over the different methods for the two data sets, based on the test set. The standard deviation is also reported.
Supervised Elliptic IBM-AML
Methods AUC-ROC AUC-PR AUC-ROC AUC-PR
Intrinsic features $0.877\pm 0.008$ $0.587\pm 0.022$ $0.755\pm 0.005$ $0.019\pm 0.001$
Egonet features $0.877\pm 0.008$ $0.604\pm 0.021$ $0.753\pm 0.006$ $0.018\pm 0.001$
DeepWalk $0.833\pm 0.009$ $0.562\pm 0.022$ $0.758\pm 0.005$ $\boldsymbol{0.019\pm 0.001}$
Node2vec $0.838\pm 0.010$ $0.622\pm 0.020$ $\boldsymbol{0.759\pm 0.005}$ $\boldsymbol{0.019\pm 0.001}$
GCN $0.862\pm 0.009$ $0.615\pm 0.020$ $0.572\pm 0.002$ $0.012\pm 0.000$
GraphSAGE $0.865\pm 0.008$ $0.637\pm 0.020$ $0.574\pm 0.002$ $0.011\pm 0.000$
GAT $\boldsymbol{0.892\pm 0.007}$ $\boldsymbol{0.679\pm 0.017}$ $0.267\pm 0.007$ $0.008\pm 0.000$
GIN $0.833\pm 0.010$ $0.551\pm 0.019$ $0.745\pm 0.006$ $0.018\pm 0.001$
Table 13: Threshold-dependent metrics: Precision, recall and F1-score values for the top $1\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
Elliptic Supervised
Methods Precision Recall F1-score
Intrinsic features $0.874\pm 0.032$ $0.155\pm 0.006$ $0.263\pm 0.010$
Egonet features $0.928\pm 0.025$ $0.164\pm 0.005$ $0.279\pm 0.008$
DeepWalk $0.870\pm 0.034$ $0.154\pm 0.006$ $0.262\pm 0.010$
Node2vec $0.973\pm 0.016$ $0.172\pm 0.003$ $0.292\pm 0.005$
GCN $\boldsymbol{1.000\pm 0.000}$ $\boldsymbol{0.177\pm 0.007}$ $\boldsymbol{0.301\pm 0.010}$
GraphSAGE $0.981\pm 0.014$ $0.174\pm 0.006$ $0.295\pm 0.009$
GAT $\boldsymbol{1.000\pm 0.000}$ $0.176\pm 0.006$ $0.299\pm 0.009$
GIN $0.977\pm 0.019$ $0.172\pm 0.006$ $0.293\pm 0.009$
Table 14: Threshold-dependent metrics: Precision, recall and F1-score values for the top $1\%$ scores over the different methods for the IBM-AML data set, based on the test set. The standard deviation is also reported.
IBM-AML Supervised
Methods Precision Recall F1-score
Intrinsic features $0.021\pm 0.005$ $0.024\pm 0.005$ $0.022\pm 0.005$
Egonet features $0.020\pm 0.005$ $0.023\pm 0.005$ $0.021\pm 0.00$
DeepWalk $0.0209\pm 0.0044$ $0.0231\pm 0.0049$ $0.0220\pm 0.0046$
Node2vec $\boldsymbol{0.0214\pm 0.0050}$ $0.0237\pm 0.0055$ $\boldsymbol{0.0225\pm 0.0052}$
GCN $0.0010\pm 0.0010$ $0.0012\pm 0.0011$ $0.0011\pm 0.0011$
GraphSAGE $0.0094\pm 0.0003$ $\boldsymbol{1.0000\pm 0.0000}$ $0.0186\pm 0.0006$
GAT $0.0011\pm 0.0011$ $0.0012\pm 0.0012$ $0.0012\pm 0.0011$
GIN $0.0210\pm 0.0040$ $0.0231\pm 0.0042$ $0.0220\pm 0.0041$
For the threshold-independent metrics, we focus on the AUC-PR, since this is more appropriate for high class imbalance. For both data sets, the models with intrinsic and those with egonet features give already good results. The suspicious transactions can be detected quite well using these “simpler” features.
For the Elliptic data set, we see that the GNNs give a slightly worse performance in terms of AUC-ROC, except for the GAT, but mostly outperform the other models when looking at the more informative AUC-PR. The difference become apparent when looking at the threshold-dependent metrics in Tables 13E1E3 and E5. Precision, recall and F1-scores are remarkably higher among the observations with highest money laundering propensity.
One surprising result is the underperformance of the GIN method for the threshold-independent metrics. Two explanations are possible. The first is that the neural network in the aggregation step is not optimal, and should be tuned further. This was not part of this experimental set-up because this would add additional hyperparameter for the depth and size of the layers for each of the layers of the GNN, which would increase the number of possible value combinations massively. Another explanation is that the inclusion of neural network aggregation actually makes GIN more sensitive to the high class imbalance, because much more weights need to be trained. This might indicate that this model is less suited for anti-money laundering. It should be noted that good performance is observed for the precision, recall and F1-scores, indicating that GIN might still be a good model when only interested in reducing false positives in the top-percentiles of predictions.
For the IBM dataset, with AUC-ROC in Table 12, the picture is difference. Here, the network information coming from DeepWalk and node2vec give additional information and increase the performance of the model. However, when looking at the GNNs, the performance is bad. The performance of GAT has even dropped to an AUC-ROC of $0.27$, meaning that the inverse prediction actually holds more value. The threshold-dependent metrics show a similar picture, but in Tables 14E2E4 and E6 it seems that DeepWalk gives better performance in the top-percentiles.
Another anomalous observations is GraphSAGE, with perfect recall, but low precision. It seems that GraphSAGE is finding all money laundering cases, but at the cost of having a high fall positive rate. This might be caused by GraphSAGE giving a flat percentage for each observation. We are working with relative cut-offs, so having the cut-off equal to the top-$1\%$, classifies all observations as money laundering.
One explanation for the poor performance of GNNs might be related to the network structure, as presented in Section 4.2. Given that the average path length is small, and the nodes are very close to each other in the network, the GNN embedding might suffer from over-smoothing [79]. This can result in very similar embeddings for the different nodes with insufficient discriminatory power.
Moreover, a deeper analysis reveals that the training loss, when training on the train and validation set, is highly unstable. Depending on the epoch and specific weight initialization, the results vary widely, with the models sometimes just predicting the majority class. For this specific data set, GNNs seem to face instability due to the extreme class imbalance.
The supervised learning experiments on both data sets demonstrate that network analytics can provide additional insights for anti-money laundering. However, depending on the network topology and the extent of the class imbalance, GNNs might be too unstable. Additional experiments are needed to quantify the sensitivity of GNNs to the imbalance, but this is outside the scope of this research.
5.2 Unsupervised Learning
The results of the unsupervised methods using isolation forests are given in Tables 15. None of the results on the Elliptic data set are particularly good, with AUC-ROC values below 0.5 indicating that the inverse of the scores would perform better as a classifier. For completeness, we also include the threshold-dependent metric in Table 16, where most models fail to detect money laundering cases.
Further analysis is done to understand why the results for unsupervised learning might be this bad. Figure 6 shows the anomaly scores coming from the isolation forest trained on the intrinsic features for the Elliptic data set. Here, a low (very negative) score signifies anomalies. All money laundering cases have scores near zero, within the bulk of the distribution, indicating that the isolation forest does not recognize these cases as anomalies.
Conversely, the results for IBM-AML are more in line with expectations. Network information captured by DeepWalk, results in the best performing model here. Comparing these results to those in Table 12, we see that DeepWalk even outperforms the supervised learning methods in terms of AUC-PR.
Figure 6: The anomaly scores for the legitimate and money laundering cases. More negative is more anomalous.
Table 15: Threshold-independent metrics: AUC-ROC and AUC-PR values over the different methods for the two data set, based on the test set. The standard deviation is also reported.
Unsupervised Elliptic IBM-AML
Methods AUC-ROC AUC-PR AUC-ROC AUC-PR
Intrinsic features $0.134\pm 0.007$ $0.058\pm 0.001$ $0.710\pm 0.015$ $0.017\pm 0.003$
Egonet features $0.217\pm 0.010$ $0.013\pm 0.000$ $0.700\pm 0.033$ $0.011\pm 0.005$
DeepWalk $0.182\pm 0.008$ $0.056\pm 0.001$ $\boldsymbol{0.744\pm 0.023}$ $\boldsymbol{0.024\pm 0.008}$
Node2vec $0.190\pm 0.008$ $0.056\pm 0.0005$ $0.632\pm 0.031$ $0.005\pm 0.001$
Table 16: Threshold-dependent metrics: Precision, recall and F1-score values for the top $1\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
Elliptic Unsupervised
Methods Precision Recall F1-score
Intrinsic features $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
Egonet features $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
DeepWalk $0.005\pm 0.005$ $0.001\pm 0.001$ $0.001\pm 0.001$
Node2vec $0.003\pm 0.004$ $000.000\pm 0$ 0.000 $0.001\pm 0.001$
Table 17: Threshold-dependent metrics: Precision, recall and F1-score values for the top $1\%$ scores over the different methods for the IBM-AML data set, based on the test set. The standard deviation is also reported.
IBM-AML Unupervised
Methods Precision Recall F1-score
Intrinsic features $0.042\pm 0.005$ $0.156\pm 0.020$ $0.067\pm 0.008$
Egonet features $0.027\pm 0.013$ $0.098\pm 0.048$ $0.042\pm 0.021$
DeepWalk $0.055\pm 0.009$ $0.204\pm 0.034$ $0.087\pm 0.015$
Node2vec $0.007\pm 0.002$ $0.025\pm 0.008$ $0.010\pm 0.003$
The most plausible explanation in our opinion lies in the assumptions underlying isolation forest and the data generating process. Two specific assumptions by Liu et al. [82]:
  • Anomalies yield fewer partitions since anomalies occupy regions with lower density;
  • Instances with distinct attributes are separated earlier in the tree, contributing to shorter paths.
The Elliptic data set is a real-life data set, where criminals actively attempt to cover their tracks [10]. By doing so, they actually become suspicious again. They try to mimic the ‘average behaviour’, but no real transaction/person is average on all attributes. In contrast, the IBM-AML data set is synthetic, with patterns and behaviours deliberately introduced. It is likely that the more complex camouflage behaviour is not present in the latter. This insight highlights the importance of knowing the exact assumptions underlying the data sets used. Synthetic data sets might give overly optimistic results, since complex behaviour inherent to the problem at hand are not captured.
6 Conclusion
In this paper, we present the results of a systematic literature review of the literature on network analytics for anti-money laundering, and of an experimental evaluation of the state-of-the-art. The systematic literature review resulted in a selection of 97 papers. These papers were processed in two steps to achieve (1) a comprehensive understanding of the broader literature and (2) detailed insight into research that has been well received by the community. The comprehensive analysis classified all papers according to seven categories, each having different sub-characteristics. The detailed analysis was performed on the $10\%$ most-cited papers in scope, which were further categorized using the framework recently proposed by Bockel-Rickermann et al. [14]. Information was given on the different processing steps taken, the challenges faced or addressed in the papers, and the definition of nodes and edges in the networks.
We concluded that a large part of the literature concerns methods that rely heavily on expert-knowledge by applying manual feature engineering and rule-based methods. Additionally, most methods find patterns that are indicative for money laundering by relying on centrality metrics and other manually engineered features. However, most methods are specifically tailored to solving just a small part of the money laundering puzzle. This might indicate a latent consensus that one should not aim at developing a single, overarching model for detecting money laundering. Instead, one may benefit from operating multiple and complementary models in parallel, each detecting different money laundering patterns. This preference appears to be driven, in part, by the lack of interpretability in more complex network representation methods.
Our analysis identified five critical gaps that future research must address:
  1. 1.
    Interpretability: Few studies have explored tools to explain the outputs of graph neural networks (GNNs), which is essential for regulatory and operational transparency.
  2. 2.
    Unsupervised Methods: Limited exploration of classic unsupervised methods restricts their potential to detect anomalies, a key requirement as money laundering tactics evolve.
  3. 3.
    Dynamic Network Analysis: Most approaches rely on static networks, with insufficient research into dynamic methods that could better capture the temporal nature of transactions.
  4. 4.
    Incomplete Labels: The inherent incompleteness of AML labels, due to undetected cases and resource constraints, highlights the need for specialized PU-learning techniques.
  5. 5.
    Prioritized Predictions: Limited resources necessitate actionable, prioritized predictions, which could be achieved through learning-to-rank methods.
Additionally, for regular fraud analytics, cost-sensitive learning is an established set of methods [14, 134, 62, 107], while none of the papers covered in our work use the cost of misclassification in the learning of their methods. Although estimation of misclassification costs in money laundering is much harder than in transaction/credit card fraud, we believe it is a branch of the literature that is worth-wile exploring more.
To extend upon the existing body of knowledge, we implemented an extensive experimental evaluation covering manual feature engineering, shallow and deep representation learning, both in a supervised and unsupervised setting. From the supervised learning results, we can conclude that network features bring additional predictive power to AML models. Depending on the setting, GNNs perform better than the manual features.
In the unsupervised learning experiments, we employed isolation forests and observed notable differences between the real-world Elliptic dataset and the synthetic IBM-AML dataset. Although on synthetic data, the methods performed well, the observations from the Elliptic data seemed to exhibit strong camouflaging tactics. The anomaly scores of the latter were so low that investigating the least suspicious transactions for money laundering became the better strategy. We conclude from this that care should be taken when testing methods on synthetic data, since it can give overly optimistic results. Future work should also investigate if this aberration still holds for other unsupervised learning methods.
For network visualisation, future work should focus on quantifying the usefulness of visualisations in helping investigators detect new patterns and money laundering cases. This can be extended into a comparative study that looks at the strengths and weaknesses of different visualisations. This requires qualitative empirical work studying how people use the visualisations, not just how it is intended to be used.
There are also some limitations in our experimental set-up that can be addressed in future work. First, conclusions are drawn based on only two data sets. To generalise our conclusions, additional tests need to be performed on a broader range of data sets. This can be done by including more open-source data sets when these become available, and rely on practitioners to use our code to test the methods and publish results on their proprietary data. Second, the decoders were always based on neural networks. The application and impact analysis of applying different machine learning methods as down-stream classifiers is another interesting area of study to extend the experiments presented here. Third, the hyperparameter tuning process was limited. While Optuna provided an efficient method for hyperparameter search, the 50 to 100 tuning rounds used in this sdy may be insufficient. The choices in this paper were made to keep everything tractable and feasible given the limited resources available. Future work could expand the parameter ranges and number of tuning rounds using more computing power.
Acknowledgments
This work was supported by the Research Foundation – Flanders (FWO research project 1SHEN24N) and by the BNP Paribas Fortis Chair in Fraud Analytics. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government.
Appendix A Literature Overview
Table A1: Literature classification. Classification of the papers according to different categories: publication data, learning method, evaluation metric, objective, and data.
Publication Data Evaluation Metric Objective Data Network
Reference Year Review Paper Bitcoin/Crypto Accuracy Precision Recall F1 TPR FPR AUC-ROC AUC-PR Time Client Classification Transaction Classification Community Detection Flow/Chain Detection Proprietary Open-Source Synthetic Inverse/undirected network Multiple networks
Dreżewski et al. [33] 2015
Ngai et al. [100] 2011
Malm and Bichler [88] 2013
Fronzetti Colladon and Remondi [41] 2017
Zhou et al. [156] 2017
Didimo et al. [29] 2011
McGinn et al. [89] 2016
Senator et al. [116] 1995
Wu et al. [142] 2021
Gao and Ye [42] 2007
Joaristi et al. [69] 2018
Irwin and Turner [64] 2018
Michalak and Korczak [94] 2011
Wu et al. [143] 2021
Li et al. [80] 2020
Phetsouvanh et al. [106] 2018
Zhou et al. [159] 2017
Didimo et al. [31] 2019
Joaristi et al. [70] 2019
Camino et al. [19] 2017
Velasco et al. [136] 2021
Nandhini and Das [98] 2016
Zhdanova et al. [155] 2014
Imanpour et al. [63] 2019
Turner et al. [128] 2019
Ovelgönne et al. [102] 2012
Magalingam et al. [85] 2015
Helmy et al. [59] 2016
Xueshuo et al. [150] 2021
Plaksiy et al. [108] 2018
Garcia-Bedoya et al. [44] 2021
Mehmet and Wijesekera [91] 2013
Xia et al. [145] 2021
Kute et al. [75] 2021
Mahootiha et al. [87] 2021
Shaikh et al. [118] 2021
Turner et al. [129] 2020
Ricci [112] 2020
Bahulkar et al. [11] 2018
Chmielewski and Stkapor [24] 2017
Zhou et al. [157] 2022
Gasanova et al. [46] 2018
Magomedov et al. [86] 2018
Tharani et al. [125] 2021
Jin and Qu [68] 2018
Mehmet et al. [92] 2014
Xiao et al. [146] 2018
Mohan et al. [96] 2022
Alarab et al. [4] 2020
Shen et al. [119] 2021
Prado-Romero and Gago-Alonso [109] 2017
Zhang et al. [154] 2021
Garcia-Bernardo et al. [45] 2022
He et al. [58] 2021
Micale et al. [93] 2019
Bhalerao et al. [13] 2019
Omar et al. [101] 2014
Winiecki et al. [141] 2020
Starnini et al. [121] 2021
Hamid [54] 2017
Gulino et al. [52] 2021
Adriaens et al. [1] 2019
Chang et al. [22] 2007
Kurshan and Shen [74] 2020
Jin et al. [66] 2022
Bellomarini et al. [12] 2021
Sousa Lima et al. [120] 2022
Çağlayan and Bahti̇yar [17] 2022
Jin et al. [67] 2022
Kurniawan et al. [73] 2015
Sun et al. [122] 2022
Gerbrands et al. [47] 2022
Dumitrescu et al. [34] 2022
Yu et al. [153] 2022
García and Mateos [43] 2021
Wang et al. [138] 2022
Gong et al. [48] 2022
Li et al. [81] 2022
Li et al. [78] 2022
Assumpção et al. [8] 2022
Semenov et al. [115] 2017
Didimo et al. [30] 2012
Cardoso et al. [20] 2022
Day [27] 2021
Alarab et al. [5] 2020
Badalov et al. [9] 2018
Korczak and Łuszczyk [72] 2011
Lokanan [84] 2022
Rathore et al. [111] 2022
Liu et al. [83] 2012
Xiong and Zhong [148] 2020
Shaikh and Nazir [117] 2018
Yang et al. [151] 2022
Xiong and Xiao [147] 2021
Mehmet and Wijesekera [90] 2013
Cheong and Si [23] 2010
Han et al. [56] 2021
Table A2: Literature classification. Classification of the papers according to the different method building blocks.
Publication Data Method Type Modelling Method
Reference Year Bitcoin/Crypto Supervised Unsupervised Semi-Supervised Mixed Visualisation Rule-Based Manual Features Walk-based Shallow Repr. Deep Repr. Correlation-Based Tree-Based SVM-Based Neural Networks Anomaly Detection Clustering Logistic Regression
Dreżewski et al. [33] 2015
Malm and Bichler [88] 2013
Fronzetti Colladon and Remondi [41] 2017
Zhou et al. [156] 2017
Didimo et al. [29] 2011
McGinn et al. [89] 2016
Senator et al. [116] 1995
Wu et al. [142] 2021
Joaristi et al. [69] 2018
Michalak and Korczak [94] 2011
Wu et al. [143] 2021
Li et al. [80] 2020
Phetsouvanh et al. [106] 2018
Zhou et al. [159] 2017
Didimo et al. [31] 2019
Joaristi et al. [70] 2019
Camino et al. [19] 2017
Velasco et al. [136] 2021
Nandhini and Das [98] 2016
Zhdanova et al. [155] 2014
Imanpour et al. [63] 2019
Turner et al. [128] 2019
Ovelgönne et al. [102] 2012
Magalingam et al. [85] 2015
Helmy et al. [59] 2016
Xueshuo et al. [150] 2021
Plaksiy et al. [108] 2018
Garcia-Bedoya et al. [44] 2021
Mehmet and Wijesekera [91] 2013
Xia et al. [145] 2021
Mahootiha et al. [87] 2021
Shaikh et al. [118] 2021
Turner et al. [129] 2020
Ricci [112] 2020
Bahulkar et al. [11] 2018
Chmielewski and Stkapor [24] 2017
Zhou et al. [157] 2022
Gasanova et al. [46] 2018
Magomedov et al. [86] 2018
Tharani et al. [125] 2021
Jin and Qu [68] 2018
Mehmet et al. [92] 2014
Xiao et al. [146] 2018
Mohan et al. [96] 2022
Shen et al. [119] 2021
Prado-Romero and Gago-Alonso [109] 2017
Zhang et al. [154] 2021
Garcia-Bernardo et al. [45] 2022
He et al. [58] 2021
Micale et al. [93] 2019
Bhalerao et al. [13] 2019
Omar et al. [101] 2014
Winiecki et al. [141] 2020
Starnini et al. [121] 2021
Hamid [54] 2017
Gulino et al. [52] 2021
Adriaens et al. [1] 2019
Chang et al. [22] 2007
Jin et al. [66] 2022
Bellomarini et al. [12] 2021
Sousa Lima et al. [120] 2022
Çağlayan and Bahti̇yar [17] 2022
Jin et al. [67] 2022
Kurniawan et al. [73] 2015
Sun et al. [122] 2022
Dumitrescu et al. [34] 2022
Yu et al. [153] 2022
García and Mateos [43] 2021
Wang et al. [138] 2022
Gong et al. [48] 2022
Li et al. [81] 2022
Li et al. [78] 2022
Assumpção et al. [8] 2022
Didimo et al. [30] 2012
Cardoso et al. [20] 2022
Alarab et al. [5] 2020
Badalov et al. [9] 2018
Korczak and Łuszczyk [72] 2011
Rathore et al. [111] 2022
Liu et al. [83] 2012
Xiong and Zhong [148] 2020
Shaikh and Nazir [117] 2018
Yang et al. [151] 2022
Xiong and Xiao [147] 2021
Mehmet and Wijesekera [90] 2013
Cheong and Si [23] 2010
Appendix B Analysis of Categories of Full Scope of Papers
In this part of the appendix, the figures are given that summarise the full scope of this literature review.
(a) The distribution of the learning methods.
(b) The evolution of the learning methods over the years.
Figure B1: The learning methods.
(a) The distribution of the model building blocks.
(b) The evolution of the model building blocks over the years. Since there are many modelling methods, the graphs are split.
Figure B2: The building blocks of the model.
(a) The distribution of the objective.
(b) The evolution of the objective over the years.
Figure B3: The objective of the method.
(a) The distribution of the evaluation metrics.
(b) The evolution of the evaluation metrics over the years. Since there are many metrics, the graphs are split.
Figure B4: The evaluation metrics.
(a) The distribution of the nature of the data.
(b) The evolution of the nature of the data over the years.
Figure B5: The nature of the data.
Appendix C Analysis of Categories between Crypto and Non-Crypto Papers
This part of the appendix gives the plots in which we compare the difference in nature of the crypto-literature compared non-crypto-related research.
Figure C1: The data for papers dealing with crypto currencies (1) and those that do not (0).
(a) The evolution of the data over the years for papers not covering crypto currencies.
(b) The evolution of the data over the years for papers covering crypto currencies.
Figure C2: Evolution of the data.
Figure C3: The distribution of the objective of the papers.
Figure C4: The distribution of the evaluation metrics used.
Appendix D The Top-Cited Papers
In this part of the appendix, we give the figures that summarise the top-cited papers.
Figure D1: The methods for the top-cited papers.
Figure D2: The modelling methods for the top-cited papers.
Figure D3: The performance metrics for the top-cited papers.
Figure D4: The objective of the top-cited papers.
Figure D5: The nature of the data of the top-cited papers.
Appendix E Results Threshold-Dependent Metrics
In this part of the appendix, we give the precision, recall and F1-scores for the models using different thresholds. The thresholds are set such that the observations with the top $0.1\%$, $10\%$ and $p\%$ of scores are classified as money laundering, where $p\%$ is equal to the relative occurrence of the labels.
Table E1: Threshold-dependent metrics: Precision, recall and F1-score values for the top $0.1\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
Methods Precision Recall F1-score
Intrinsic features $0.061\pm 0.117$ $0.001\pm 0.002$ $0.002\pm 0.005$
Egonet features $0.589\pm 0.186$ $0.012\pm 0.004$ $0.023\pm 0.007$
DeepWalk $0.637\pm 0.147$ $0.013\pm 0.003$ $0.025\pm 0.006$
Node2vec $0.984\pm 0.048$ $0.019\pm 0.002$ $0.038\pm 0.003$
GCN $\boldsymbol{1.000\pm 0.000}$ $0.019\pm 0.001$ $0.037\pm 0.001$
GraphSAGE $0.942\pm 0.073$ $\boldsymbol{0.020\pm 0.003}$ $\boldsymbol{0.039\pm 0.005}$
GAT $\boldsymbol{1.000\pm 0.000}$ $0.019\pm 0.001$ $0.037\pm 0.001$
GIN $\boldsymbol{1.000\pm 0.000}$ $0.019\pm 0.001$ $0.038\pm 0.002$
Table E2: Threshold-dependent metrics: Precision, recall and F1-score values for the top $0.1\%$ scores over the different methods for the IBM-AML data set, based on the test set. The standard deviation is also reported.
Methods Precision Recall F1-score
Intrinsic features $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
Egonet features $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
DeepWalk $\boldsymbol{0.018\pm 0.012}$ $0.002\pm 0.002$ $0.004\pm 0.003$
Node2vec $0.007\pm 0.009$ $0.001\pm 0.001$ $0.002\pm 0.002$
GCN $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
GraphSAGE $0.009\pm 0.000$ $\boldsymbol{1.000\pm 0.000}$ $\boldsymbol{0.019\pm 0.001}$
GAT $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
GIN $0.010\pm 0.010$ $0.001\pm 0.001$ $0.002\pm 0.002$
Table E3: Threshold-dependent metrics: Precision, recall and F1-score values for the top $10\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
Methods Precision Recall F1-score
Intrinsic features $0.357\pm 0.011$ $0.629\pm 0.019$ $0.456\pm 0.014$
Egonet features $0.361\pm 0.010$ $0.636\pm 0.018$ $0.460\pm 0.013$
DeepWalk $0.347\pm 0.011$ $0.611\pm 0.019$ $0.443\pm 0.014$
Node2vec $0.351\pm 0.011$ $0.617\pm 0.019$ $0.447\pm 0.014$
GCN $0.348\pm 0.019$ $0.614\pm 0.021$ $0.444\pm 0.020$
GraphSAGE $0.358\pm 0.016$ $0.633\pm 0.018$ $0.458\pm 0.016$
GAT $\boldsymbol{0.379\pm 0.015}$ $\boldsymbol{0.665\pm 0.018}$ $\boldsymbol{0.482\pm 0.015}$
GIN $0.341\pm 0.016$ $0.600\pm 0.019$ $0.435\pm 0.016$
Table E4: Threshold-dependent metrics: Precision, recall and F1-score values for the top $10\%$ scores over the different methods for the IBM-AML data set, based on the test set. The standard deviation is also reported.
Methods Precision Recall F1-score
Intrinsic features $0.015\pm 0.001$ $0.161\pm 0.013$ $0.027\pm 0.002$
Egonet features $0.014\pm 0.001$ $0.157\pm 0.012$ $0.026\pm 0.002$
DeepWalk $\boldsymbol{0.015\pm 0.001}$ $0.165\pm 0.012$ $\boldsymbol{0.028\pm 0.002}$
Node2vec $0.015\pm 0.001$ $0.164\pm 0.013$ $0.027\pm 0.002$
GCN $0.012\pm 0.000$ $0.986\pm 0.004$ $0.023\pm 0.001$
GraphSAGE $0.009\pm 0.000$ $\boldsymbol{1.000\pm 0.000}$ $0.019\pm 0.001$
GAT $0.002\pm 0.000$ $0.017\pm 0.004$ $0.003\pm 0.001$
GIN $0.015\pm 0.001$ $0.160\pm 0.012$ $0.027\pm 0.002$
Table E5: Threshold-dependent metrics: Precision, recall and F1-score values for the top $p=2\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
Methods Precision Recall F1-score
Intrinsic features $0.327\pm 0.010$ $0.635\pm 0.019$ $0.432\pm 0.013$
Egonet features $0.334\pm 0.009$ $0.650\pm 0.018$ $0.441\pm 0.012$
DeepWalk $0.316\pm 0.010$ $0.614\pm 0.019$ $0.417\pm 0.013$
Node2vec $0.318\pm 0.010$ $0.619\pm 0.019$ $0.420\pm 0.013$
GCN $0.317\pm 0.017$ $0.618\pm 0.020$ $0.419\pm 0.019$
GraphSAGE $0.329\pm 0.015$ $0.641\pm 0.018$ $0.435\pm 0.015$
GAT $\boldsymbol{0.350\pm 0.014}$ $\boldsymbol{0.677\pm 0.018}$ $\boldsymbol{0.461\pm 0.015}$
GIN $0.315\pm 0.015$ $0.612\pm 0.020$ $0.416\pm 0.016$
Table E6: Threshold-dependent metrics: Precision, recall and F1-score values for the top $p=0.11\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
Methods Precision Recall F1-score
Intrinsic features $0.001\pm 0.004$ $0.000\pm 0.001$ $0.000\pm 0.001$
Egonet features $0.001\pm 0.003$ $0.000\pm 0.001$ $0.000\pm 0.001$
DeepWalk $\boldsymbol{0.017\pm 0.012}$ $0.002\pm 0.002$ $0.004\pm 0.003$
Node2vec $0.011\pm 0.009$ $0.002\pm 0.001$ $0.003\pm 0.003$
GCN $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
GraphSAGE $0.009\pm 0.000$ $\boldsymbol{1.000\pm 0.000}$ $\boldsymbol{0.019\pm 0.001}$
GAT $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
GIN $0.009\pm 0.009$ $0.001\pm 0.001$ $0.002\pm 0.002$